OpenStudy (anonymous):

Use a half-angle identity to find the evact value of each function. A.) Tan 165 degrees I have tried and cannot come up with the right answer. Can I have some help?

OpenStudy (anonymous):

Exact not evact sorry

OpenStudy (tkhunny):

2*165º = 330º = 360º - 30º Reference Angle -30º

OpenStudy (anonymous):

\[\tan \left( \frac{x}{2} \right) = \frac{\sin x}{1 + \cos x}\] plug x = 330º

OpenStudy (anonymous):

i did that, and i checked the answer in the back and it did not go right. I messed up somewhere but i do not know where.

OpenStudy (anonymous):

maybe you forgot to rationalize the denominator?

OpenStudy (tkhunny):

Did you consider the quadrant? 330º is Quadrant IV (Tangent is Negative). 165º is Quadrant II (Tangent is negative) Please show your work and we'll be able to see where you wander off.

OpenStudy (anonymous):